OpenStudy (tinash):

Solve this simultaneously:

OpenStudy (tinash):

2b + c=6 b +c=6

OpenStudy (tinash):

change the sign of the second one.

OpenStudy (yumyum247):

3b + 2c = 12

OpenStudy (yumyum247):

water we exactly solving for thow?

OpenStudy (tinash):

Basically I am find the general term of a quadratic Formula: Tn=an^2 + bn + c 2a=4 a=2 T1=8 2(1)^2 + b(1) + c=8 2+b+c=8 -------------- b+c=6 T2=14 2(2)^2+b(2)+c=14 8+2b+c=14 -------------- 2b+tc=6

OpenStudy (yumyum247):

@knov

OpenStudy (knov):